(a-5)^2+(b-c)^2+(c~d)^2+(b+c+d-9)^2=0 then (a+b+c)(b+c+b) ?

Grade:12

1 Answers

Vijay Mukati
8 years ago
Dear Student,

IF given equation is equal to zero and since each term is squared term, therefore each term should be equal to zero,
i.e. a-5 = 0, b – c = 0, c – d = 0 and b+c+d-9 = 0.
Using these, find the value of a, b, c, and d and then finally find the value of desired expression.

Thanks.